OpenStudy (anonymous):

Solve 3 log base 5 y - 1og base 25 y = 10

OpenStudy (aravindg):

well first try to use property of logarithm on first log term

OpenStudy (anonymous):

do u know how to start?

OpenStudy (anonymous):

no i didnt attend the lesson.

OpenStudy (aravindg):

ie. \[\large \log x^n=n \log x\]

OpenStudy (anonymous):

so u dont know any properties of logarithm?

OpenStudy (anonymous):

i copied down all the 5 formulas without knowing the explanation as i was absent from class.

OpenStudy (aravindg):

well by missing your class , there is no use tring to solve any problem unless u hav worked on theory part yourself

OpenStudy (aravindg):

i think it will be a waste of time for both you and us

OpenStudy (anonymous):

so u know that, \[\log_{n ^{m}} a =1/m \log_{n} a\]

OpenStudy (anonymous):

could you solve it for me as i try to figure out.

OpenStudy (aravindg):

I think a better idea would be trying to learn the basics first and then attempt the question

OpenStudy (anonymous):

I didnt miss the basic lesson.

OpenStudy (aravindg):

i just found a good document for you to read and understand http://www.shoreline.edu/math/Handouts/Properties%20of%20Logarithms.pdf

OpenStudy (anonymous):

@Sgstudent actuallly @AravindG is right...first understand the properties then u can solve the problem quite easily!

OpenStudy (aravindg):

make a good reading of that ,when you become confident about the basic properties let us know ! we will be happy to help you :)

OpenStudy (anonymous):

okay thnx

OpenStudy (aravindg):

infact if you do the basics thoroughly you can do this question in your head ! imagine that ! :)

OpenStudy (anonymous):

Still, could you at least help me in this question? What are the steps involved?

OpenStudy (aravindg):

we will once you make an effort to learn the basics

OpenStudy (aravindg):

you can run only after you learn to walk!

OpenStudy (anonymous):

I said i learnt the basics already but i am stuck with the question.

OpenStudy (aravindg):

why are you still after the question? it would only take half an hour to read through that document and familiarize with properties of logarithm .Please make an effort !

OpenStudy (aravindg):

infact that document contains practise problem+answers!!

OpenStudy (anonymous):

I read thrice already.

OpenStudy (anonymous):

It's okay if you dont believe me , peace out.

OpenStudy (aravindg):

did you do the practise problems ?

OpenStudy (anonymous):

AravindG